# Rethinking Invariance Regularization in Adversarial Training to Improve Robustness-Accuracy Trade-off

### v3

@@ 3.4 Auto-blance for reducing hyperparameters
where $\alpha_{(t)}$ and $\beta_{(t)}$ are the hyperparameters at the current epoch $t$, and $Acc_{(t-1)}$ is the clean accuracy of the model $f_{\theta}^{\text{auxBN}}$ at the previous epoch $(t-1)$. This adjustment automatically forces the model to focus more on adversarial images as its accuracy on clean images improves. We empirically demonstrate that this heuristic dynamic adjustment strategy works well in practice, successfully reducing the hyperparameter tuning cost. We provide an ablation study in Sec. 5.5
@@ 5.1 Analysis and Understanding of each component of AR-AT
Figure 3: Mixture distribution problem, indicated by the L2 distance between adversarial and clean feature, $||z-z^{\prime}||_{2}$. While $\mathcal{L}_{V0}$ already suffers from this problem, the use of stop-grad exacerbates this issue by weakening invariance regularization.
@@
To understand the mixture distribution problem, we analyze the similarity between adversarial and clean features. In Fig. 3, we plot the L2 distance between the adversarial and clean features, $||z-z^{\prime}||_{2}$, for the same input $x$. The increasing distance over time indicates that adversarial and clean features become more dissimilar during training, illustrating a robustness-accuracy trade-off. This suggests that shared BN layers between clean and adversarial inputs may fail to accurately estimate adversarial feature statistics (i.e., mixture distribution problem). Appendix F.2 demonstrates that the estimated batch statistics become more stable with the use of split-BN.
@@
(b) The gradient norm of the loss function ($||\nabla_{\theta}L||_{2}$).
@@
Latent projection with predictor MLP head enhances performance by improving training stability. Finally, we analyze the effectiveness of the predictor MLP head. Tab. 2 shows that employing the predictor MLP head ($\mathcal{L}_{V1}$ vs. $\mathcal{L}_{V2}$) leads to further improvements in the robustness and accuracy. In Fig. 4, we analyze the training stability. Fig. 4a shows the cosine distance between the adversarial representations at epochs $t$ and $t+1$, where the predictor MLP stabilizes the updates of the representations where invariance loss is applied. This suggests that the predictor MLP head helps stabilize the optimization of the invariance regularization by smoothing the updates of the representations. Fig. 4b reveals that, indeed, the predictor MLP reduces the gradient norm of the loss function ($||\nabla_{\theta}L||_{2}$), implying more stable optimization.
@@ 5.5 Ablation study
ResNet Penultimate layer (“layer4") 86.28 51.78
@@
WRN Penultimate layer (“layer3") 88.87 52.50
@@ B.5 Ablation Study: Contributions of Each Component in AR-AT
(4) $\mathcal{L}_{V2}$checkmark $\checkmark$ 83.00 49.14 0.63
@@
(8) $\mathcal{L}_{V3}$checkmark $\checkmark$ $\checkmark$ 86.29 52.40 0.52
@@ F.1 Further analysis of $L2\;norm(z^{\prime}-z)$
In Figure 13, we show the L2 distance between adversarial and clean features ($||z^{\prime}-z||_{2}$) for ResNet-18 trained on CIFAR-10. In addition to Figure 3 from the main text, we include results for standard training and adversarial training (AT) without invariance regularization for comparison.
All methods show increasing in $||z^{\prime}-z||_{2}$ over time, reflecting the robustness-accuracy trade-off. As the model accuracy improves, adversarial perturbations become easier to find, leading to higher $||z^{\prime}-z||_{2}$. This poses a challenge for methods using both clean and adversarial inputs during training: the diverging distributions of clean and adversarial samples make it difficult for BatchNorm (BN) layers to correctly estimate batch statistics, leading to the mixture distribution problem. Note that standard training and AT avoid this issue by using only clean or adversarial samples, respectively.
Stop-grad increases $||z^{\prime}-z||_{2}$ compared to AR-AT without it. While stop-grad resolves gradient conflict, it weakens the invariance regularization by removing the second term in Eq. 5 ($(Dist(z^{\prime},sg(z))+Dist(z,sg(z^{\prime})))/2$). This leads to an increase in $||z^{\prime}-z||_{2}$, exacerbating the mixture distribution problem. Our split-BN strategy effectively mitigates this issue, enhancing the utility of stop-grad.
Figure 13: L2 distance between adversarial and clean features ($||z^{\prime}-z||_{2}$). All training methods show an increase in $||z^{\prime}-z||_{2}$ over time, reflecting the robustness-accuracy trade-off. Using stop-grad increases $||z^{\prime}-z||_{2}$ compared to AR-AT without stop-grad, exacerbating the mixture distribution problem.

### v4

@@ 3.4 Auto-blance for reducing hyperparameters
where $\alpha_{(t)}$ and $\beta_{(t)}$ are the hyperparameters at the current epoch $t$, and $Acc_{(t-1)}$ is the clean accuracy of the model $f_{\theta}^{\text{auxBN}}$ at the previous epoch (t1)(t-1). This adjustment automatically forces the model to focus more on adversarial images as its accuracy on clean images improves. We empirically demonstrate that this heuristic dynamic adjustment strategy works well in practice, successfully reducing the hyperparameter tuning cost. We provide an ablation study in Sec. 5.5
@@ 5.1 Analysis and Understanding of each component of AR-AT
Figure 3: Mixture distribution problem, indicated by the L2 distance between adversarial and clean feature, $z-z^{\prime}||_{2}$. While $\mathcal{L}_{V0}$ already suffers from this problem, the use of stop-grad exacerbates this issue by weakening invariance regularization.
@@
To understand the mixture distribution problem, we analyze the similarity between adversarial and clean features. In Fig. 3, we plot the L2 distance between the adversarial and clean features, $z-z^{\prime}||_{2}$, for the same input $x$. The increasing distance over time indicates that adversarial and clean features become more dissimilar during training, illustrating a robustness-accuracy trade-off. This suggests that shared BN layers between clean and adversarial inputs may fail to accurately estimate adversarial feature statistics (i.e., mixture distribution problem). Appendix F.2 demonstrates that the estimated batch statistics become more stable with the use of split-BN.
@@
(b) The gradient norm of the loss function ($\nabla_{\theta}L||_{2}$).
@@
Latent projection with predictor MLP head enhances performance by improving training stability. Finally, we analyze the effectiveness of the predictor MLP head. Tab. 2 shows that employing the predictor MLP head ($\mathcal{L}_{V1}$ vs. $\mathcal{L}_{V2}$) leads to further improvements in the robustness and accuracy. In Fig. 4, we analyze the training stability. Fig. 4a shows the cosine distance between the adversarial representations at epochs $t$ and $t+1$, where the predictor MLP stabilizes the updates of the representations where invariance loss is applied. This suggests that the predictor MLP head helps stabilize the optimization of the invariance regularization by smoothing the updates of the representations. Fig. 4b reveals that, indeed, the predictor MLP reduces the gradient norm of the loss function ($\nabla_{\theta}L||_{2}$), implying more stable optimization.
@@ 5.5 Ablation study
ResNet
Penultimate layer (“layer4") 86.28 51.78
@@
WRN
Penultimate layer (“layer3") 88.87 52.50
@@ B.5 Ablation Study: Contributions of Each Component in AR-AT
(4) $\mathcal{L}_{V2}$ $\checkmark$ $\checkmark$ 83.00 49.14 0.63
@@
(8) $\mathcal{L}_{V3}$ $\checkmark$ $\checkmark$ $\checkmark$ 86.29 52.40 0.52
@@ F.1 Further analysis of $L2\;norm(z^{\prime}-z)$
In Figure 13, we show the L2 distance between adversarial and clean features ($z^{\prime}-z||_{2}$) for ResNet-18 trained on CIFAR-10. In addition to Figure 3 from the main text, we include results for standard training and adversarial training (AT) without invariance regularization for comparison.
All methods show increasing in $z^{\prime}-z||_{2}$ over time, reflecting the robustness-accuracy trade-off. As the model accuracy improves, adversarial perturbations become easier to find, leading to higher $z^{\prime}-z||_{2}$. This poses a challenge for methods using both clean and adversarial inputs during training: the diverging distributions of clean and adversarial samples make it difficult for BatchNorm (BN) layers to correctly estimate batch statistics, leading to the mixture distribution problem. Note that standard training and AT avoid this issue by using only clean or adversarial samples, respectively.
Stop-grad increases $z^{\prime}-z||_{2}$ compared to AR-AT without it. While stop-grad resolves gradient conflict, it weakens the invariance regularization by removing the second term in Eq. 5 ($(Dist(z^{\prime},sg(z))+Dist(z,sg(z^{\prime})))/2$). This leads to an increase in $z^{\prime}-z||_{2}$, exacerbating the mixture distribution problem. Our split-BN strategy effectively mitigates this issue, enhancing the utility of stop-grad.
Figure 13: L2 distance between adversarial and clean features ($z^{\prime}-z||_{2}$). All training methods show an increase in $z^{\prime}-z||_{2}$ over time, reflecting the robustness-accuracy trade-off. Using stop-grad increases $z^{\prime}-z||_{2}$ compared to AR-AT without stop-grad, exacerbating the mixture distribution problem.